OpenStudy (anonymous):

how do i find the value of SIN 12 CSC 12 without a calculator

OpenStudy (anonymous):

wait, is it sin(12)*csc(12) ?

jimthompson5910 (jim_thompson5910):

Hint: csc(x) = 1/sin(x)

OpenStudy (anonymous):

yes it is multiplied @macknojia

OpenStudy (anonymous):

sin(12)*(1/sin(12))= sin(12)/sin(12)= ?

OpenStudy (anonymous):

=1?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

what about sec^2 233- tan^2 233=

jimthompson5910 (jim_thompson5910):

check out these identities http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

are you aware of the property sin^2x+sin^2x=1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin^2x+cos^2x=1 ?**** divide both sides by cos^2x and you'll get the identity you are trying to find.

OpenStudy (anonymous):

instead of an x you have the angle in your problem which is 233.

OpenStudy (anonymous):

yes it equals 1

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):



OpenStudy (anonymous):

the top "B" should be "C"